OpenStudy (anonymous):

Find the sum of the first 50 terms of the sequence below. an=3n+2

OpenStudy (anonymous):

@sourwing help me out please?

OpenStudy (anonymous):

a1 = 3.1 +2 a2 = 3.2 +2 a3= 3.3 +2 a4= 3.4 +2 a5........... ....................... ...................... a50=3.50 +2 adding all we have required sum = 3(1+2+3+4+..........50) +2*50 =3(50)(50+1)/2 +100 =3*25*51+100 =3925